The areas of the bases of a frustum of a regular pyramid is 250 m2 and 640 m2 respectively. Find the altitude of the pyramid if the altitude of its frustum is 130 m?
1
Expert's answer
2020-12-13T19:19:38-0500

Let "A_1,A_2" and "h_1" be the areas of upper base,lower base and height of the frustum.

Also let "H" be the height of the pyramid.

Then,Volume of pyramid is "=(1\/3)\u00d7A_2\u00d7H"

and Volume of frustum is "=(1/3)×h_1×[A_1+A_2+\sqrt{(A_1.A_2)}]"

We know that,

Volume of pyramid "=" Volume of frustum "+"Volume of pyramid containing upper base of frustum as its base. ..........(1)

Let "h_2" be the height of the pyramid containing upper base of frustum as its base.

Then we have , "H=h_1+h_2"

"\implies h_2=H-h_1" .......(2)

Here we have given, "A_1=250" "m^2" , "A_2=640" "m^2" , "h_1=130" "m"

From (2) we have, "h_2=H-130"

From (1) we have,

"(1/3)×A_2×H=[(1/3)×h_1×[A_1+A_2+\sqrt {(A_1.A_2)}]]+[(1/3)×A_1×h_2]"

Therefore,

"(1/3)×640×H=[(1/3)×130×[640+250+\sqrt {(640×250)}]]+[(1/3)×(H-130)×250]"

"\implies 640H-250H=130(1290-250)"

"\implies 390H=130×1040"

"\implies H=(1040/3)"

Therefor the required height of the pyramid is

"=(1040\/3)" "m"